OpenStudy (anonymous):

Find the angle between the vectors v=2i+2j+k and w=2i+2i-k

OpenStudy (jamesj):

I'm guessing you mean w = 2i + 2j - k. Then as the dot product/scalar product \[ v.w = |v| |w| \cos \theta \] Solve this equation for \( \cos \theta \) and then for \( \theta \).

OpenStudy (anonymous):

\[\text{ArcCos }\left[\frac{7}{9}\right]=0.679674=38.94\text{ degrees} \]

OpenStudy (anonymous):

correct thanks!